To investigate how often families camp in the summer, The Outdoorsman magazine surveyed 1,100 adults living with children under the age of 18. The survey results are shown in the table below.


1
Expert's answer
2021-02-04T14:19:53-0500

The survey results are shown in the table below.  

"\begin{matrix}
 \text{\# of times in summer} & \text{\# of survey resoinses} \\
 0 & 100 \\
1 & 120 \\
2 & 180 \\
3 & 200 \\
4 & 280 \\
5 & 100 \\
6 & 80 \\
\geq7 & 40 \\
\end{matrix}"

"N=1100"


a) If "A" is the event that a family does not camp out, then


"P(A)=\dfrac{100}{1100}=\dfrac{1}{11}\approx0.091"

b) If "B" is the event that a family camps at least 3 times during the summer, then


"P(B)=\dfrac{200+280+100+80+40}{1100}=\dfrac{7}{11}\approx 0.64"



c) If "C" is the event that a family camps at most 2 times during the summer, then

"P(C)=\dfrac{100+120+180}{1100}=\dfrac{4}{11}\approx 0.36"